OpenStudy (anonymous):

I dont think I am doing this right... Find the average rate of change of y with respect to x from P to Q. 1) y = x3 - 7; P(5, 118), Q(5.1, 125.651)

OpenStudy (anonymous):

use the formula for the slope \[\frac{125.651-118}{5.1-5}\]

OpenStudy (anonymous):

hmmm... is that " y = \[x ^{3} -7\] " ?

OpenStudy (anonymous):

I didn't get that but I think I entered it incorrectly

OpenStudy (anonymous):

what did you get as a number?

OpenStudy (anonymous):

\(125.651-118=7.651\) then divide by \(.1\) same as multiply by 10 you get \(76.51\)

OpenStudy (anonymous):

I wasn't dividing by .1 I left it at 7.651

OpenStudy (anonymous):

ah that was the mistake

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

hmmm... I didnt get the part where you wrote the equation "y = x3-7"... what was the point of that ?.. well, if you got it, thats fine then ^^